OpenStudy (anonymous):

If 72.5 grams of calcium metal (Ca) react with 65.0 grams of oxygen gas (O2) in a synthesis reaction, how many grams of the excess reactant will be left over when the reaction is complete? Be sure to write out the balanced equation for this reaction and to show all of your work. @Nurali

OpenStudy (anonymous):

Can you start this off yourself by writing the balanced equation?

OpenStudy (anonymous):

I dunno how @ghuczek

OpenStudy (anonymous):

It is a synthesis, or combination reaction. Does that help to figure out the equation for the chemical reaction?

OpenStudy (anonymous):

Sorry, I have to go. Class is starting ...

OpenStudy (anonymous):

Oh, okay thanks anyway! @ghuczek

OpenStudy (nurali):

2Ca + O2 -----> 2CaO 2 mol Ca reacts with 1 mol O2 72.5 g Ca x (1 mol / 40.08 g) = 1.81 mol Ca 65.0 g O2 x (1mol / 32.0 g) = 2.03 mol O2 1.81 mol Ca x (1 mol O2 / 2 mol Ca) = 0.905 mol O2 (needed to use up all the Ca) Excess mol O2 = 2.03 - 0.905 = 1.12 mol O2 left over 1.12 mol O2 x (32.0 g / 1 mol) = 36.0 g O2 left over

OpenStudy (anonymous):

THANK YOU SIR!! @Nurali

OpenStudy (nurali):

My Pleasure.